# BatchLLM: Optimizing Large Batched LLM Inference with Global Prefix Sharing and Throughput-oriented Token Batching

### v1

@@ 6.2.1. Microbenchmark Evaluation
We synthesize the dataset for the evaluation according to the data distribution of the industry workload, where the average share-degree (i.e., the number of the distinct prompt that share the same prefix) is about 7, the average prefix length is about 1100 tokens, and the average distinct prompt length is about 400 tokens. We sample 3000 queries for this case study.
We conduct this experiment on the NVIDIA A100 GPU, using Mistral-7B model. Table.1 shows the effectiveness of BatchLLM of this workload (and another industry workload that will describe later), with about 1.3$\times$ speedup over the best configuration of vLLM. BatchLLM better reuses the KV context than the vLLM baseline with the optimization in Sec.4.2, for which we have the breakdown analysis in Sec.6.3.1. BatchLLM also better mixes the decoding tokens with prefill tokens to increase the size of token-batches with the optimization in Sec.4.3, for which we have the breakdown analyze in Sec.6.3.2. It also benefits from the optimized Attention kernel described in this paper.
6.2.3. Case Study 2: Short Prefix with Long Distinct Prompt.
We have evaluated another industry workload with different data distribution to that in Sec.6.2.2. This is a ranking job that ranks a set of web documents given the user query according to the relevance. Specifically, if forms the query-document pairs where the same query along with the instruction is the shared prefix across a number of documents. The difference to the workload in Sec.6.2.2 is that, the length of the common prefix is shorter than that of the distinct prompt (i.e., document). However, the sharing degree in this case is relatively large. We evaluate this workload with 1000 samples, where the average share-degree is 11.36, the length of the common prefix is about 770, and the length of the distinct prompt is about 978.
We conduct this experiment on the AMD MI200 GPU, using Qwen2-7B model. BatchLLM still gets a significant gain comparing with all the other vLLM settings as shown in Table.1, 1.47×1.47$\times$1.47 × speedup than the best performance of vLLM. Even though the shared prefix has a shorter length, the large sharing degree still makes it achieve a high token saving ratio, which is about 38%. As a result, it still benefits from the explicit common prefix reusing.
@@ 6.3.1. Reusing Effect Comparison
We compare the saving ratio (refer to Eq.2) between vLLM’s prefix-caching and BatchLLM on the industry task in Sec.6.2.2. The saving ratio of vLLM’s prefix-caching is 44.2%, and BatchLLM achieves 54.9% for this metric, showing that the explicit prefix identification can help to reuse more KV context. Note that we only use 3000 requests to evaluate the saving ratio. With the increase of the requests number, it can be expected that the implicit LRU cache policy can suffer more from the eviction of the reusable KV context.
@@ 6.3.2. Token-batching Analysis
Figure 6. The token number at each iteration with the token-batching optimization in Sec.4.3, for the same workload with that in Fig.2. It significant reduces the "vallyes" compared to the baseline in Fig.2. Note that the prefix-sharing is not enable here, showing the efficacy to the general scenarios of batched/offline inference.
To investigate the transferability of token-batching optimization in Sec.4.3, we conducted ablation experiments on the memory-centric token-batching and group-based reordering/scheduling, as shown in Fig.7. Meanwhile, we propose a metric suitable for Large Batched LLM data processing, namely the end-to-end serving time excluding the cool-down (TECD). This metric excludes the cool-down phase of the service (as shown in the rightmost part of Fig.6), because as the data scale increases, the time of the cool-down phase remains constant (it is related to the number of iterations of the few remaining requests in the engine). By profiling TECD, it can better reflect the actual processing time of ultra-large-scale data with the same length distribution from a batch of smaller-scale data.
(a) The throughput comparison.
(b) The normalized time comparison excluding the cool-down (TECD).
Figure 7. The performance comparison between enabling token-batching or reordering and the baseline. The bar of "None" is the vLLM 0.5.4 baseline with or without BatchLLM’s global prefix-sharing optimization. The left side is the vanilla vLLM and the right side is with BatchLLM’s prefix-sharing.
Fig.7a shows the throughput improvement of enabling group reordering (Sec.4.3.2), memory-centric token batching (Sec.4.3.3), and the both. We conducted experiments using Mistral-7B on the task described in Sec.6.2.2 (with 3000 requests), and achieved a throughput improvement of up to 9%. The left side is the baseline that does not enabling the prefix-sharing. This is to demonstrate that this optimization is general and can be applied to the non-shared workloads. The right side is for the prefix-sharing scenario. It shows that each of the two techniques can achieve better performance than the valina strategy (the "None" bar in the figure). And the two optimization can work together well for the prefix sharing scenarios.
According to Fig.7b, by applying strategies of BatchLLM token-batching, we can reduce the TECD time by up to 15%, and to some extent compress the overall number of steps excluding the cool-down (about 1600 -> about 1350) while keeping the batch size almost unchanged.
6.3.3. Kernel Performance Comparison
@@
Figure 8. The performance comparison between the baseline, CascadeInfer and BatchLLM. For one chunked-prefill batch or one decoding batch, the setting i/j/k (like 256/2048/all) means the length of global shared prefix, the length of distinct parts for each request, and how many requests with one single shared prefix in one single prefix-sharing group.
We compare BatchLLM’s prefix-shared Attention implementation with several baselines on both NVIDIA A100 GPU and AMD MI200 GPU. Fig.8 shows the performance comparison of different implementations, including Triton official FlashAttention implementation [^9] https://github.com/triton-lang/kernels/blob/main/kernels/flash_attention.py, official FlashAttention [^10] https://github.com/Dao-AILab/flash-attention, FlashinferCascade (Ye et al., 2024a), and BatchLLM’s kernels. We compare the performance among different kernels under two scenarios: the pure decoding scenario (request count 256) and the chunked-prefill scenario where a token-batch includes both prefill chunks and decoding tokens (7 prefill requests and 256 decoding requests). Besides that, how many requests one single prefix-group have, and various length setting of both shared prefix and non-shared tokens, are non-trivial. Our experiments are conducted based on these insights.
@@ 6.3.4. Global Prefix Identification Overhead
We have measured the time of the global prefix identification described in Sec.4.2. The time range of the overhead starts before adding the token ids of each prompt to form the basic prefix tree and ends after generating the final list of prefix-sharing groups. As for the industry workload in Sec.6.2.2, the total overhead of the 3000 requests is 0.28 seconds. While the time to process the 3000 requests with the input of prefix-sharing groups format is 233.46 seconds. The global prefix identification procedure nearly has no overhead compared with the end-to-end execution time.
@@ 7. Related Work
Attention kernel optimization. Memory-efficient Attention (Rabe and Staats, 2021) and FlashAttention (Dao et al., 2022) fuse the Attention operators into a single kernel to boost the performance, with Online Softmax (Milakov and Gimelshein, 2018) to tile the Attention computation. The recent prefix-shared Attention optimization (ChunkAttention (Ye et al., 2024b), RelayAttention (Zhu et al., 2024), Hydragen (Juravsky et al., 2024), and Cascade Inference (Ye et al., 2024a)) compute the attention on the prefix and other part separately in different kernels, for which the former is transformed from matrix-vector multiplication between Q and K/V into MatMul, and reduce the results of different parts according to the Online Softmax algorithm. Different from these works, BatchLLM fuses the Attention computation of different parts into the same kernel to reduce tail latency and kernel launch overhead. The horizontal fusion has been proposed and used in the existing machine learning optimizers (Pan et al., 2023; Li et al., 2022). BatchLLM borrows this idea and applies it in the prefix-shared Attention.
@@ 8. Conclusion
This paper presents BatchLLM, the holistic optimization techniques for large batched LLM-based information processing and management tasks. It identifies the limitations of existing methods, and optimizes the tasks with the global information of the large batch. It explicitly extracts the common prefix globally to avoid prefix’s early eviction problem, and simplifies the prefix pattern by enlarging the first-level prefix with a DP algorithm on the tree to reduce the overhead of scheduling and Attention computation. It schedules the requests at the granularity of prefix-sharing groups, which enables the global prefix sharing the best and shrinks the lifetime of prefix’s KV memory. It proposes the request reordering and memory-centric token batching method to better mix the prefill chunks into the decoding token-batches and thus better saturates the GPU. Finally, it presents the horizontal fused prefix-shared Attention kernel to reduce the tail effect and kernel launch overhead. Extensive evaluation shows that BatchLLM outperforms vLLM by $\times1$ to 2.0×2.0$\times$2.0 × on a set of microbenchmarks and two typical industry workloads.

### v2

@@ 6.2.2. Industry Workload Evaluation
We synthesize the dataset for the evaluation according to the data distribution of the industry workload (Fig.7), where the average share-degree (i.e., the number of the distinct prompt that share the same prefix) is about 3, the average common prefix length is about 1570 tokens, and the average distinct prompt length is about 30 tokens. We sample 8000 queries for this case study.
We conduct this experiment on the NVIDIA A100 GPU and AMD MI200 GPU, using Qwen-2.5-7B model. Table.1 shows the effectiveness of BatchLLM of this workload (and another industry workload that will describe later), with about 1.3$\times$/1.26$\times$ speedup over the best configuration of vLLM. BatchLLM better reuses the KV context than the vLLM baseline with the optimization in Sec.4.2, for which we have the breakdown analysis in Sec.6.3.1. BatchLLM also better mixes the decoding tokens with prefill tokens to increase the size of token-batches with the optimization in Sec.4.3, for which we have the breakdown analyze in Sec.6.3.2. It also benefits from the optimized Attention kernel described in this paper.
@@ 6.3.1. Reusing Effect Comparison
To show that BatchLLM optimizes the KV reuse, we evaluate the throughput and the token saving ratio in both microbenchmarks and the industry task.
Figure 8. The token saving (or reusing) ratio of different sharing-degrees and different shared-prefix lengths on A100, based on the microbenchmark. This evaluation is based on Llama-3.1-8B-Instruct, while the length of non-shared context is 200, and the generated length is 100.
As shown in Fig.8, we can see the token saving ratio is improved in BatchLLM across different settings. And it becomes higher when the sharing-degree of requests is larger. Specifically, in the microbenchmark setting where length of shared prefix is 2000 and the sharing-degree is 16, BatchLLM achieves 85.2% token saving ratio, the upper bound in this scenario, comparing to the 40.1% token saving ratio in vLLM + chunked-prefill + prefix-caching. For the industry workload, BatchLLM also makes a token saving ratio of 58.1%, while the number of vLLM + chunked-prefill + prefix-caching is 35.8%. It is worth mentioning that BatchLLM can achieve the best token reusing since it performs explicit common prefix management. With the increase of the requests number and the length of prefix, it can be expected that the implicit LRU cache policy of vLLM’s prefix-caching can suffer more from the eviction of the reusable KV context, while the inference could benefit more from BatchLLM. For example, when the length of shared prefix is 16000, and the share degree is 16, the token reusing ratio of vLLM and SGLang are 6.3% and 5.2%, while the reusing ratio of BatchLLM is 92.6 %, much better than the other two baselines.
@@ 6.3.2. Token-batching Analysis
Figure 9. Token number per iteration with token-batching optimization (Sec.4.3), showing significant valley reduction compared to Fig.2’s baseline. Prefix-sharing disabled to demonstrate general batched/offline inference efficacy.
We evaluated our token-batching optimization through ablation studies on throughput-oriented token-batching. Testing on Qwen-2.5-7B with 3000 industry workload query (Fig.2) shows effective performance improvement. Results demonstrate effectiveness both with and without prefix-sharing, with combined optimizations working particularly well in prefix-sharing scenarios.
Table 2. Token-Batching Breakdown.
Settings Disable token-batching Enable token-batching
vLLM + c 5.79 5.94
BatchLLM 8.16 8.47
@@
Figure 10. The performance comparison between the baseline, CascadeInfer and BatchLLM. The setting i/j/k (like 256/2048/all) means the length of global shared prefix, the length of distinct parts for each request, and how many requests with one single shared prefix in one single prefix-sharing group.
6.3.3. Kernel Performance Comparison
We compare BatchLLM’s prefix-shared Attention implementation with several baselines on both NVIDIA A100 GPU and AMD MI200 GPU. Fig.10 shows the performance comparison of different implementations, including Triton official FlashAttention implementation [^10] https://github.com/triton-lang/kernels/blob/main/kernels/flash_attention.py, official FlashAttention [^11] https://github.com/Dao-AILab/flash-attention, v2.6.1, FlashinferCascade [^12] https://github.com/flashinfer-ai/flashinfer, v0.1.4 (Ye et al., 2024a), and BatchLLM’s kernels. We compare the performance among different kernels under two scenarios: the pure decoding scenario (request count 256) and the chunked-prefill scenario where a token-batch includes both prefill chunks and decoding tokens (7 prefill requests and 256 decoding requests). Besides that, how many requests one single prefix-group have, and various length setting of both shared prefix and non-shared tokens, are non-trivial. Our experiments are conducted based on these insights.
@@ 6.3.4. Global Prefix Identification Overhead
We have measured the time of the global prefix identification described in Sec.4.2. The time range of the overhead starts before adding the token ids of each prompt to form the basic prefix tree and ends after generating the final list of prefix-sharing groups. As for the industry workload in Sec.6.2.2, the total overhead of the 8000 requests is 1.51 seconds. While the time to process the 8000 requests with the input of prefix-sharing groups format is 919.54 seconds. The global prefix identification procedure nearly has no overhead compared with the end-to-end execution time.
@@ 7. Related Work
Attention kernel optimization. Memory-efficient Attention (Rabe and Staats, 2021) and FlashAttention (Dao et al., 2022) fuse the Attention operators into a single kernel to boost the performance, with Online Softmax (Milakov and Gimelshein, 2018) to tile the Attention computation. The recent prefix-shared Attention optimization (ChunkAttention (Ye et al., 2024b), RelayAttention (Zhu et al., 2024), Hydragen (Juravsky et al., 2024), and Cascade Inference (Ye et al., 2024a)) compute the attention on the prefix and other part separately in different kernels, for which the former is transformed from matrix-vector multiplication between Q and K/V into MatMul, and reduce the results of different parts according to the Online Softmax algorithm. Different from these works, BatchLLM fuses the Attention computation of different parts into the same kernel to reduce tail latency and kernel launch overhead. The horizontal fusion has been proposed and used in the existing machine learning optimizers (Pan et al., 2023, 2024; Li et al., 2022). BatchLLM borrows this idea and applies it in the prefix-shared Attention.
@@ 8. Conclusion
This paper presents BatchLLM, the holistic optimization techniques for large batched LLM-based information processing and management tasks. It identifies the limitations of existing methods, and optimizes the tasks with the global information of the large batch. It explicitly extracts the common prefix globally to avoid prefix’s early eviction problem, and simplifies the prefix pattern by enlarging the first-level prefix with a DP algorithm on the tree to reduce the overhead of scheduling and Attention computation. It schedules the requests at the granularity of prefix-sharing groups, which enables the global prefix sharing the best and shrinks the lifetime of prefix’s KV memory. It proposes the request reordering and memory-centric token batching method to better mix the prefill chunks into the decoding token-batches and thus better saturates the GPU. Finally, it presents the horizontal fused prefix-shared Attention kernel to reduce the tail effect and kernel launch overhead. Extensive evaluation shows that BatchLLM outperforms vLLM and SGLang by 1.3×1.3$\times$1.3 × to 10.8×10.8$\times$10.8 × on a set of microbenchmarks and a typical industry workload under different hardware environments.